Alexander2003:

Kylie recorded the number of pears on a pear tree at the end of every week. The following function represents the number of pears on the pear tree, where x represents the number of weeks. P(x) = 2^(2 + x) Which statement is true? The number of pears on the tree quadruples each week. The number of pears on the tree increases by 2 each week. The number of pears on the tree increases by 4 each week. The number of pears on the tree doubles each week. mod note: Please do not give away the answer directly. Low-effort and/or plagiarized answers will be removed.

jhonyy9:

where are the statements here ?

Vocaloid:

Missing information has been added.

jhonyy9:

P(x)= 2^(2+x) 2^(2+x) = (2^2) *(2^x) = 4 * 2^x hope this will help you